Find an equation in the slope-intercept form for the line that passes through (2,3) with a slope of -3/8. y = __x + ___.
Thank you for your help.

Answer by Alan3354(69443)

Find an equation in the slope-intercept form for the line that passes through (2,3) with a slope of -3/8
----
y-y1 = m*(x-x1)
y-3 = (-3/8)*(x-2)